4.Samantha drove 440 km in 5 hours. She drove part of the way at 80 km/h and the rest at 100 km/h. What distance did she drive at each speed? (5 marks)
Answer by jorel555(1290)

If x and y are the hours she drove at the different rates, then:
x+y=5
80x+100y=440
-------
80x+100y=440
80x+80y=400
20y=40
y=2
x=3
She drove 3 hours at 80 kmh and 2 hours at 100kmh!!!!!!!